OpenStudy (anonymous):

What is the variance of the following data? If necessary, round your answer to two decimal places. 600, 470, 170, 430, 300

OpenStudy (kirbykirby):

The mean, is \[\Large \bar{x}=\frac{600+ 470+ 170+ 430+ 300}{5}=394 \] Assuming this is the sample variance, then \[\large \frac{1}{n-1}\sum_{i=1}^n(x_i-\bar{x})^2\\ ~ \\\ =\frac{1}{5-1}\left((600-\bar{x})^2+(470-\bar{x})^2+(170-\bar{x})^2+(430-\bar{x})^2+(300-\bar{x})^2 \right) \] where you replace \(\bar{x}\) with 394